Six added to 3 times a number is equal to 4 less than 5 times the number. What is the number?




A.
10

B.
5

C.
–5

D.
–10

Answer by ewatrrr(24785)

 
Hi,
Six added to 3 times a number "x" is equal to 4 less than 5 times the number.
3x + 6 = 5x-4 ****
10 = 2x
5 = x
CHECKING our Answer***
15 + 6 = 25 - 4 = 21